# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ V Execution time of the QSP algorithm
Figure 4(a) shows a comparison of the excitation probability after applying the BB1 sequence. The IBM-Q result is shown by orange dots. Qubit dynamics simulations performed with QuTiP [41, 42, 43] (including relaxation and dephasing), for signals defined by AIM using both direct QSP analogy and double LZSM transitions are shown by the red curve. All results show good agreement with the expected polynomial distribution shown by the blue curve.
@@
In Fig. 4(b) we compare the time needed to perform the BB1 sequence by different methods. Note that for the angle $\theta=0$ in the QSP algorithm, with direct QSP to AIM analogy, time tends to infinity, as discussed before. This divergence arises due to the definition of the driving frequency in Eq. (21) and the probability of the LZSM transition Eq. (14c), which yields $\omega=0$, so the period of the driving tends to infinity in this case. Note also that the BB1 execution time by the double LZSM transitions results in a stable operation time for all angles $\theta$.
@@
We found that the IBM-Q, cloud based quantum computer, needed significantly more time to execute the BB1 sequence, especially for the non-optimized version, where the dependence on the number of gates appears. This is mainly due to the Rabi-like resonance pulse type of the driving used in current quantum computers, which loses fidelity significantly with faster gates. Despite the much faster gate execution time of the optimized algorithm with maximum optimization on an IBM-Q (where the sequence of all single-qubit gates is replaced by just one gate), the double LZSM transition is still several times faster. But the LZSM transitions require a small energy gap in the anti-crossing point and the distance between levels should grow fast with detuning $\varepsilon$. Therefore, the speed advantage of the LZSM-based gates may be constrained by hardware limitations in practical implementations.
VI Conclusions
@@
Quantum signal processing is a powerful tool for implementing polynomial transformations and other quantum algorithms with high precision. Due to its structural similarity to the adiabatic-impulse model, it is possible to establish a direct mapping between the parameters of these two approaches, enabling efficient translation of one algorithm into the other. This analogy provides a novel perspective for the direct implementation of QSP in quantum systems.
@@
Furthermore, the use of adiabatic-impulse model, with double Landau-Zener-Stückelberg-Majorana transitions, demonstrates significantly reduced time for implementing quantum signal processing algorithm. The double LZSM transitions allow us to solve the infinite transition time problem for $\mathcal{P}=0$ in the direct implementation of QSP with AIM.
@@
The double LZSM transitions approach is a full analog of the Mach-Zehnder interferometer, which allows to use the double transition approach from AIM in a programmable quantum gate for quantum photonic computations [38, 39, 40], so the same approaches can be used for both of them. With the double LZSM transition, we show how to perform rotations around the $x$-axis on the Bloch sphere for any desired angle $\theta$. By combining this with idling times to perform signal processing operators, we can use fast nonadiabatic LZSM gates to perform any QSP sequence and achieving faster execution time even for the most optimized case for the IBM-Q, where all single-qubit gates are combined into one. In principle, the same single-qubit gate sequence combination technique can be applied for LZSM gates to reduce the execution time accordingly.
@@
The adiabatic-impulse model describes well the dynamics of systems with small energy gaps and can be used as an alternative approach to implement fast quantum logic gates and algorithms such as QSP.
@@ ACKNOWLEDGMENTS
S.N.S. acknowledges support via the Virtual Short-Term Scientist Engagement Fellowship Program for Ukrainian Scientists with Dual-Use Relevant Expertise (STCU, grant # NSEP1030). F.N. is supported in part by: the Japan Science and Technology Agency (JST) [via the CREST Quantum Frontiers program Grant No. JPMJCR24I2, the Quantum Leap Flagship Program (Q-LEAP), and the Moonshot R&D Grant Number JPMJMS2061], and the Office of Naval Research (ONR) Global (via Grant No. N62909-23-1-2074).

### v3

@@ V Execution time of the QSP algorithm
Figure 4(a) shows a comparison of the excitation probability after applying the BB1 sequence. The IBM-Q result is shown by orange dots. Qubit dynamics simulations were performed with QuTiP [47, 48, 49], including relaxation and dephasing with similar rates to IBM-Q, for signals defined by AIM using both direct QSP analogy and double LZSM transitions are shown by the red curve. All results show good agreement with the expected polynomial distribution shown by the blue curve.
@@
In Fig. 4(b), we compare the time needed to perform the BB1 sequence by different methods. Note that for the angle $\theta=0$ in the QSP technique, with direct QSP to AIM analogy, time diverges, as discussed before. This divergence arises due to the definition of the driving frequency in Eq. (21) and the probability of the LZSM transition Eq. (14f), which yields $\omega=0$, so the period of the driving tends to infinity in this case. Note also that the BB1 execution time by the double LZSM transitions results in a stable operation time for all angles $\theta$.
@@
We found that the IBM-Q cloud-based quantum computer needed significantly more time to execute the BB1 sequence, especially for the non-optimized version, where the dependence on the number of gates appears. This is mainly due to the Rabi-like resonance pulse type of the driving used in current quantum processors, which loses fidelity significantly with faster gates. Despite the much faster gate execution time of the optimized algorithm with transpilation on an IBM-Q (the sequence of all single-qubit gates is replaced by one gate), the double LZSM transitions are still several times faster. But the LZSM transitions require a small energy gap in the anti-crossing point, and the distance between levels should grow fast with detuning $\varepsilon$. Therefore, the speed advantage of the LZSM-based gates may be constrained by hardware limitations in practical implementations.
VI Scalability to High-Dimensional Hilbert Spaces
To perform real-world quantum calculations dynamics of multi-level systems is crucial. While single-qubit dynamics is relatively simple, adding more levels increases the difficulty of the problem as $2^{N}$ for $N$ levels. In this section, we present two ways to expand our theory to multi-level systems. The first one consists in using the qubitization technique to decompose a multi-level system into many two-level systems with QSP-like evolution in each one. The second one is to directly apply the generalized AIM to a multi-level system. Both procedures are schematically presented in Fig. 5.
Figure 5: Schematic diagrams for two ways to apply the adiabatic-impulse model for multi-level systems. Panel (a) shows the qubitization technique, which allows decomposing a multi-level system by K=N(N1)/2K=N(N-1)/2 single-qubit evolution algorithms. The decomposed evolution operator can be realized with the QSP sequences, and then fast LZSM gates can be applied to perform QSP. Panel (b) demonstrates a way to directly realize multi-level evolution with generalized AIM, which consists of consecutive LZSM transitions with respective energy levels. Dashed lines represent diabatic energy levels in the two-qubit system; the orange line represents one of the possible trajectories for the occupation probability with transitions between levels on each anti-crossing point.
VI.1 From quantum signal processing to quantum singular value transformation
The algorithm we discussed in Sec. III can also be used for large qubit registers or multi-level systems via decomposing the multi-level problem to a certain number of single-qubit problems with the procedure known as the “qubitization” technique [9, 50]. While QSP is often introduced in the setting of a single-qubit subspace, the same idea extends naturally to higher-dimensional systems. For a block-encoded operator acting on many qubits, the qubitization decomposes the global Hilbert space into a direct sum of two-dimensional invariant subspaces, each labeled by a singular value. For details see Appendix A. To describe such a system, one has to
  • construct a block-encoding of the operator;
  • apply the qubitization to identify the collection of effective qubit subspaces;
  • implement the QSP phase sequence simultaneously across all of them.
In this way, QSP and its generalization QSVT perform polynomial transformations on the full operator spectrum, with complexity determined by the polynomial degree rather than the physical dimension of the system. This makes the framework equally applicable to large qubit registers or multi-level (qudit) systems. And can also be used to simulate an arbitrary quantum Hamiltonian with QSP sequences on quantum processors, since these usually have a lot of constraints in the parameters of the driving signal that does not allow to apply an arbitrary Hamiltonian directly.
VI.2 Generalized AIM
Quantum logic gates based on LZSM transitions, as described in Sec. II.2 for single-qubit operations, can also be applied to multi-qubit operations [51] with several approximations and constraints. For this, one has to
  • locally consider each anti-crossing on the energy levels time evolution as a two-level system;
  • apply the transition matrix that corresponds to the two respective levels for each anti-crossing;
  • describe other levels’ dynamics as the adiabatic evolution.
This is schematically described in Fig. 5(b). The main limitation of this approach consists in that the distance between levels should increase enough to finalize the transition process before the next transition starts. If several anti-crossings happen simultaneously between different pairs of levels, they are treated as two transitions between respective levels. The main challenge of this approach is to ensure that the relative phase between all the levels corresponds to a desired multi-level quantum gate.
VII Conclusions
@@
Quantum signal processing is a powerful tool for implementing polynomial transformations and other quantum algorithms with high precision. Due to its structural similarity to the adiabatic-impulse model, it is possible to establish a direct mapping between the parameters of these two approaches, enabling efficient translation of one algorithmic technique into the other. This analogy provides a novel perspective for the direct implementation of QSP in quantum systems.
@@
Furthermore, the use of the adiabatic-impulse model, with double Landau-Zener-Stückelberg-Majorana transitions, demonstrates significantly reduced time for implementing the quantum signal processing technique. The double LZSM transitions allow us to solve the infinite transition time problem for $\mathcal{P}=0$ in the direct implementation of QSP with AIM.
@@
The double LZSM transitions approach is a full analog of the Mach-Zehnder interferometer, which allows to use the double transition approach from AIM in a controllable quantum gate for quantum photonic computations, so the same approaches can be used for both of them. With the double LZSM transition, we show how to perform rotations around the $x$-axis on the Bloch sphere for any desired angle $\theta$. By combining this with idling times to perform signal processing operators, we can use fast non-adiabatic LZSM gates to perform any QSP sequence and achieve faster execution time even for the most optimized case for the IBM-Q, where all single-qubit gates are combined into one. In principle, the same single-qubit gate sequence combination technique can be applied for LZSM gates to reduce the execution time accordingly.
@@
The adiabatic-impulse model describes well the dynamics of systems with small energy gaps and can be used as an alternative approach to implement fast quantum logic gates and algorithmic techniques such as QSP.
@@ ACKNOWLEDGMENTS
The authors acknowledge the useful discussions with Z. M. Rossi. S.N.S. acknowledges support via the Virtual Short-Term Scientist Engagement Fellowship Program for Ukrainian Scientists with Dual-Use Relevant Expertise (STCU, Grant No. NSEP1030). F.N. is supported in part by the Japan Science and Technology Agency (JST) [via the CREST Quantum Frontiers program Grant No. JPMJCR24I2, the Quantum Leap Flagship Program (Q-LEAP), and the Moonshot R&D Grant No. JPMJMS2061], and the Office of Naval Research (ONR) Global (via Grant No. N62909-23-1-2074).
DATA AVAILABILITY
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.
Appendix A Interpretation of the quantum singular value transformation in terms of rotations
The framework of quantum signal processing (QSP) admits a natural generalization to the quantum singular value transformation (QSVT) [9]. In this setting, arbitrary matrices can be embedded into larger unitaries via block-encoding, which then allows one to implement polynomial transformations of their singular values. This embedding procedure is called qubitization, since it reduces the dynamics of a potentially large Hilbert space to effective two-level systems.
Every complex matrix $C$ admits a singular value decomposition:
$|,C=\sum_{i=1}^{r}\sigma_{i}\ket{w_{i}}\bra{v_{i}}$, (34)
[…]
The appearance of singular values is natural in block-encoding because the input and output subspaces of the encoded operator may differ in dimension. Unlike unitary operators (where eigenvalues suffice), a general matrix must be described in terms of its singular value decomposition.
Suppose we wish to encode $C$ into a larger unitary $U.$ Then there exist projectors $\Pi$ and $\widetilde{\Pi}$ such that
$C=\widetilde{\Pi}U\Pi.$ (35)
The meaning of these projectors is that $\Pi$ selects the right singular vector space (the input block of $U$) and $\widetilde{\Pi}$ selects the left singular vector space (the output block of $U$). Thus, $\Pi$ identifies where inputs are injected, and $\widetilde{\Pi}$ identifies where outputs appear. In QSVT, we also introduce phase-modified projectors:
$\Pi_{\phi}=\exp{[i\phi(2\Pi-I)]},$ (36)
$\widetilde{\Pi}_{\phi}=\exp{[i\phi(2\widetilde{\Pi}-I)]}.$ (37)
[…]
The qubitization ensures that the Hilbert space decomposes into invariant two-dimensional subspaces
[…]
with one such subspace for each singular value $\sigma_{i}$. On each $\mathcal{S}_{i}$, the block-encoded $U$ acts as an $SU(2)$ matrix,
$U|_{\mathcal{S}_{i}}=R(\sigma_{i}).$ (39)
To make unitarity explicit, we parameterize the singular value as
$\sigma_{i}=\cos\!\left(\tfrac{\theta_{i}}{2}\right),\qquad\theta_{i}=2\arccos\sigma_{i}.$
Then
$R(\sigma_{i})=R_{y}(\theta_{i})=e^{-i(\theta_{i}/2)Y}.$
Equivalently, in XXZZ language one may write
$R(\sigma_{i})=e^{i\pi Z/4}\,R_{x}(\theta_{i})\,e^{-i\pi Z/4}.$
Restricted to $\mathcal{S}_{i}$, the projector phases act diagonally:
$\Pi_{\phi}|_{\mathcal{S}_{i}}=\exp{(-i\phi Z)},\qquad\widetilde{\Pi}_{\phi}|_{\mathcal{S}_{i}}=\exp{(i\phi Z)}.$ (40)
These represent opposite $z$-axis rotations of the effective qubit.
The central building block of QSVT is the iterate [9]
$\widetilde{U}_{\boldsymbol{\Phi}}=\prod_{k=1}^{d/2}\Bigl(\Pi_{\phi_{2k-1}}\,U^{\dagger}\,\widetilde{\Pi}_{\phi_{2k}}\,U\Bigr),$ (41)
[…]
Restricted to $\mathcal{S}_{i}$, one factor of the even-$d$ product reduces to
$\Pi_{\phi_{2k-1}}U^{\dagger}\widetilde{\Pi}_{\phi_{2k}}U\big|_{\mathcal{S}_{i}}$ (42)
$\qquad=\exp{(-i\phi_{2k-1}Z)}R(\sigma_{i})\exp{(i\phi_{2k}Z)}R(\sigma_{i}).$
Here $U$ and $U^{\dagger}$ swap between right and left singular vector bases, while the exponential factors encode QSP phases. Then
$\sigma_{i}=\cos(\theta_{i}/2)$ (43)
$R(\sigma_{i})=R_{y}(\theta_{i}).$
With $ZR_{y}(\theta)=R_{y}(-\theta)Z$, the full QSVT sequence becomes
$\widetilde{U}_{\boldsymbol{\Phi}}|_{\mathcal{S}_{i}}=\prod_{k=1}^{d/2}\Bigl(R_{z}(2\phi_{2k-1})R_{y}(\theta_{i})R_{z}(2\phi_{2k})R_{y}(-\theta_{i})\Bigr),$ (44)
where $\exp{(i\phi Z)}=R_{z}(-2\phi)$. Finally, recalling that $R_{y}(\theta)$ can be rewritten using $X$-axis rotations,
$R_{y}(\theta)=\exp{\left(i\frac{\pi}{4}Z\right)}R_{x}(\theta)\exp{\left(-i\frac{\pi}{4}Z\right)},$ (45)
we obtain
$\widetilde{U}_{\boldsymbol{\Phi}}|_{\mathcal{S}_{i}}=\exp{\left(i\frac{\pi}{4}Z\right)}\Biggl[\prod_{k=1}^{d/2}$ $\Bigl(R_{z}(2\phi_{2k-1})R_{x}(\theta_{i})$ (46)
$\times$ $R_{z}(2\phi_{2k})R_{x}(-\theta_{i})\Bigr)\Biggr]\exp{\left(-i\frac{\pi}{4}Z\right)}.$
Similarly, one can get the result for the odd $d$, for which the pattern of repeatedly applied $R_{x}R_{z}$ gates is preserved.
Thus, after qubitization, each singular subspace $\mathcal{S}_{i}$ is reduced to a qubit undergoing controlled rotations about the $Z$ and $X$ axes. The sequence of phases $\{\phi_{j}\}$ fully determines the polynomial transformation implemented, while the angle $\theta_{i}$ encodes the singular value $\sigma_{i}=\cos\theta_{i}$.
In the special case where the block-encoded operator is itself unitary, all singular values have magnitude one, and QSVT reduces to the original QSP framework acting on eigenvalues lying on the unit circle. QSVT thus generalizes QSP by allowing polynomial transformations of singular values of arbitrary operators, not just phases of unitaries.